# Mathematical Language Models: A Survey

### v3

@@ 2.2.2. Theorem Proving
Ongoing research is delving into the integration of language models and interactive proof assistants within the realm of mathematics. This involves generating proofs through tactic prediction (Polu and Sutskever, 2020b; Han et al., 2022; Lample et al., 2022; Jiang et al., 2022), automating formalization processes (Wu et al., 2022; Jiang et al., 2023b), and developing unified tools (Welleck and Saha, 2023). Due to the substantial computational demands of exploration, language models applied in this domain have traditionally been limited in size. However, recent advancements have showcased potential in employing larger models (First et al., 2023; Jiang et al., 2023b). LLEMMA (Azerbayev et al., [n. d.]) demonstrates few-shot proof autoformalization and tactic prediction, offering a vast dataset of formal mathematics and an openly accessible model for further investigation in these directions. These endeavors establish crucial theoretical and practical frameworks for automating and enhancing tactic prediction.
@@ 3. PLMs-based Methods
LMs  (§3.1) [ GPT-$f$ (Polu and Sutskever, 2020a), EPT (Kim et al., 2020) , Generate &\&& Rank (Shen et al., 2021) , Thor (Jiang et al., 2022) , HTPS (Lample et al., [n. d.]) , Galactica (Taylor et al., 2022) , MATH-PLM (Hendrycks et al., 2021b) , LISA (Jiang et al., 2021)
PACT (Han et al., 2022) , Minerva (Lewkowycz et al., [n. d.]) , LIME (Wu et al., [n. d.]) , leaf, text width=54.2em ] ] [ Non-Autoregression
LMs  (§3.2) [ Aristo (Clark et al., 2021), GenBERT (Geva et al., 2020) , NF-NSM (Feng et al., 2022) , MWP-BERT (Liang et al., 2022) , MathBERT(Peng et al., 2021) , TAGOP (Zhu et al., 2021) , MT2Net (Zhao et al., 2022)
@@
Building [ Auto-explanation (NAKAMOTO et al., 2023), ProofNet (Azerbayev et al., 2023) , WizardLM (Xu et al., 2023) , Wizardmath (Luo et al., 2023) , leaf, text width=44.6em ] ] [ Instruct
Tuning [ MathGLM (Yang et al., 2023), Goat (Liu and Low, 2023), Calculon (Muffo et al., [n. d.]), PaLM 2-L-Math (Liu et al., 2023b), LLEMMA (Azerbayev et al., [n. d.]) , leaf, text width=44.6em ] ] [ In-context
Learning [ ScratchpadGPT (Nye et al., 2021) , Codex-math (Drori et al., 2022) , In-Context Sampling (Liu et al., 2022) , ComplexPrompt (Fu et al., 2023)
ART (Zhang et al., 2023c) , leaf, text width=44.6em ] ] ] [ Tool-based
@@
Methods [ SymbLLM (He-Yueya et al., 2023), PoT (Chen et al., 2022a), Codex-math (Drori et al., 2022) , MathPrompter (Imani et al., 2023) , PAL (Gao et al., 2023) , Code-scratchpad (Upadhyay et al., 2023) , leaf, text width=44.6em ] ] [ Multi-tool
Methods [ Toolformer (Schick et al., 2023), Chameleon (Lu et al., 2023a) , ART (Paranjape et al., 2023) , ToolkenGPT (Hao et al., 2023b) , CRITIC (Gou et al., 2023a) , TALM (Parisi et al., 2022)
@@
CoT [ Answer Rationales (Ling et al., 2017) , MathQA (Amini et al., 2019), CoT (Wei et al., 2022), Complexity-based CoT (Fu et al., 2023) , LVSS (Lightman et al., 2023), RFT (Yuan et al., 2023a)
Thought Propagation (Yu et al., 2023a) , AFT (Wang et al., 2023a) , PoT (Chen et al., 2022a) , AOT (Sel et al., 2023) , MAmmoTH (Yue et al., 2023) , leaf, text width=44.6em ] ] [ Construction of
CoT [ Zero-shot CoT (Kojima et al., [n. d.]) , Auto-CoT (Zhang et al., 2023c) , Complexity-based CoT (Fu et al., 2023) , PromptPG-CoT (Lu et al., 2023b)
AutoMate CoT (Shum et al., 2023) , BoostedPrompt (Pitis et al., 2023) , leaf, text width=44.6em ] ] ] [ Advanced CoT
@@
Methods [ Code-based self-verification (Zhou et al., 2023d) , VerifyCoT (Ling et al., 2023) , DIVERSE (Li et al., 2023b) , Verify-and-Edit (Zhao et al., 2023a)
Retrieval-CoT (He et al., 2022) , SCREWS (Shridhar et al., 2023) , leaf, text width=44.6em ] ] [ Ensemble-based
Methods [ Self-Consistency (Wang et al., 2023b) , Diversity-of-Thought (Naik et al., 2023) , Complexity-based CoT (Fu et al., 2023), DIVERSE (Li et al., 2023b)
@@
Methods [ ToT (Yao et al., 2023b) , TouT (Mo and Xin, 2023) , ToT-system (Long, 2023) , GoT (Yao et al., 2023a; Besta et al., 2023; Lei et al., 2023) , RESPROMPT (Jiang et al., 2023a)
RAP (Hao et al., 2023a) , LATS (Zhou et al., 2023e) , LLM+P (Liu et al., 2023a) , LLM+DP (Dagan et al., 2023) , Self-refine (Madaan et al., 2023; Sun et al., 2023) , ISR-LLM (Zhou et al., 2023b)
Reflexion (Shinn et al., 2023) , leaf, text width=44.6em ] ] [ Socratic Teaching
Methods [ Socratic Models (Zeng et al., 2022), SOCRATIC QUESTIONING (Qi et al., 2023), Socratic prompt (Chang, 2023), SocratiQ (Al-Hossami et al., 2023b)
multi-turn Socratic advice (Ang et al., 2023) , leaf, text width=44.6em ] ] ] ] ]
@@
3.1. Autoregression LMs
The distinctions between Autoregression LMs (ALMs) and Non-Autoregressive LMs (NALMs) in sequence generation tasks are depicted in Figure 3. Furthermore, ALMs can be categorized into two architectures: causal decoder (e.g., GPT-1 (Radford et al., 2018) and GPT-2 (Radford et al., 2019)) and encoder-decoder (e.g., T5 (Raffel et al., [n. d.])). The generation mechanisms of these two structures for sequences are depicted on the left and in the middle of Figure 3, respectively. These two architectures have not only significantly propelled the advancement of PLMs but have also become the predominant foundational architectures for LLMs.
To enhance the performance of PLMs in mathematical problems, many improved methods based on ALMs have been proposed. GPT-$f$ (Polu and Sutskever, 2020a) presents an automated prover and proof assistant to explore the performance of LMs to automated theorem proving. To resolve challenges related to expression fragmentation and operand-context separation, Kim et al. (Kim et al., 2020) introduce the Expression-Pointer Transformer (EPT), a pure neural model that employs both an ‘Expression’ token and operand-context pointers during the generation of solution equations. Generate &\&& Rank (Shen et al., 2021) introduces a novel ranking task for MWPs, a multi-task framework built upon a generative PLM, effectively addressing the challenge of minor mistakes in mathematical expressions. Thor (Jiang et al., 2022) introduces a framework that integrates language models and automated TP. The latter is employed to selectively choose relevant premises from an extensive library to assist the language model in TP. HTPS (Lample et al., [n. d.]) presents HyperTree Proof Search, a novel algorithm leveraging online training from past proof searches, enhancing model performance and surpassing the previous SOTA GPT-$f$. Galactica (Taylor et al., 2022) proposes a working memory token approach that can achieve strong performance over existing methods on mathematical MMLU (Hendrycks et al., 2021a) and MATH (Hendrycks et al., 2021b) benchmarks.
Furthermore, MATH-PLM (Hendrycks et al., 2021b) introduces a new dataset named MATH, consisting of challenging competition mathematics problems. It is found that relying solely on increasing budgets and model parameter counts would be impractical for achieving robust mathematical reasoning. LISA (Jiang et al., 2021) gathers an extensive collection of lemmas and theorems extracted from the Isabelle standard library and the Archive of Formal Proofs (AFP). Using this vast corpus, it builds LMs that prove theorems effectively within the AFP. PACT (Han et al., 2022) extracts abundant self-supervised data from kernel-level proof terms, enabling joint training alongside the standard tactic prediction objective. This effectively enhances the model’s ability for TP. Minerva (Lewkowycz et al., [n. d.]) trains a language model, initially pre-trained on general natural language data, through additional training on technical content to address Quantitative Reasoning Problems. It achieves a SOTA performance on technical benchmarks including MATH (Hendrycks et al., 2021b). LIME (Wu et al., [n. d.]) introduces a novel pre-training methodology that specifically learns inductive bias for mathematical reasoning. Observations indicate that models trained using LIME exhibit superior performance compared to standard transformers.
@@ 3.2. Non-Autoregression LMs
Unlike ALMs, NALMs enable the model to generate all parts of a sequence simultaneously, without depending on previously generated content, as depicted on the right side of Figure 3. Notably, models such as BERT (Kenton and Toutanova, 2019), Roberta (Liu et al., 2019), and related architectures utilize masked word representations, leveraging pre-trained context-aware embeddings as comprehensive semantic features. This significantly elevates the performance benchmarks of NLP tasks.
@@ 4. LLMs-based Methods
Large-scale language models (LLMs) are designed for processing and generating text akin to human communication (Touvron et al., 2023a; OpenAI, 2023). Mathematics is also a form of language, that communicates complex concepts and relationships through a structured system of symbols and notation, akin to the rules of a spoken language. Thus, a language model that grasps these mathematical rules can “speak” the language of mathematics, proving to be a valuable asset for mathematicians. In numerous ways, a sophisticated language model like GPT-4 (OpenAI, 2023) becomes an invaluable tool in the field of mathematics (Azerbayev et al., [n. d.]; Luo et al., 2023). We classify the existing studies into four parts: instruction learning, tool-based methods, fundamental CoT methods, and advanced CoT methods (Figure 2).
@@ 4.1. Instruction Learning
A semi-supervised approach, as presented in Auto-explanation (NAKAMOTO et al., 2023), leverages LLMs to create datasets for automating the scoring of mathematical self-explanations in the context of mathematics education. A challenging benchmark ProofNet (Azerbayev et al., 2023) is introduced to build a system of automated theorem proving. Meanwhile, prompt retrieval and distilled back translation methods are introduced for statement autoformalization. WizardLM (Xu et al., 2023) presents a groundbreaking method called Evol-Instruct, designed to autonomously generate high-quality instructions by LLMs themselves. The conceptual flow is depicted in Figure 4. The instructional process commences with a straightforward “initial instruct” and progresses into diverse forms, deepening, increase reasoning, add constraints, concretizing, complicate Input (formula), in-Breadth evolving, etc. Furthermore, Wizardmath (Luo et al., 2023) proposes a reinforced Evol-instruct method (RLEIF) to build more complexity of instruct dataset. This method effectively enhances the mathematical reasoning capabilities of Llama-2 by supervised fine-tuning and PPO training (Schulman et al., 2017).
Figure 4. Example of Evol-Instruct taken from WizardLM (Xu et al., 2023).
@@
Instruction tuning stands as a potent method to elevate the prowess of large models, aiding in steering the model towards generating outputs that align with human intent (Zhang et al., 2023a). Compared to the pre-training phase of LLMs, instruction tuning demands notably less instruction data and computational resources, making it a prevalent strategy for enhancing a model’s domain-specific capabilities. For instance, MathGLM (Yang et al., 2023) showcases that even a language model with 2 billion parameters can adeptly execute multi-digit arithmetic operations with limited training data. Goat (Liu and Low, 2023) discerns patterns in the tokenization of different numbers and proposes a consistent approach for their tokenization. Utilizing the LLaMA model, Goat ensures consistent tokenization for various numbers, illustrated in Figure 5. Similarly, Calculon (Muffo et al., [n. d.]) refines the model’s arithmetic skills by employing a digit decomposition technique to construct a fine-tuning dataset, as depicted in Figure 6.
Figure 5. Consistent tokenization for similar number in Goat (Liu and Low, 2023).
Figure 6. Number decomposition taken from Calculon (Muffo et al., [n. d.]).
In PaLM 2-L-Math (Liu et al., 2023b), three fine-tuning strategies were explored, revealing key insights: 1) The quality and style of step-by-step solutions significantly influence model performance; 2) Combining solution re-ranking and majority voting yields a more substantial performance enhancement compared to using them individually. Additionally, multi-task fine-tuning, separating solution generation and evaluation tasks, surpasses the baseline achieved by solution fine-tuning alone. Moving beyond the enhancement of mathematical capabilities through fine-tuning, LLEMMA (Azerbayev et al., [n. d.]) introduces the Proof-Pile-2 dataset, an amalgamation of mathematical texts and code. Continual pre-training with Code Llama empowers the model to leverage Python interpreters and formal theorem provers, demonstrating impressive performance on the MATH benchmark.
@@
In-context Learning (ICL) (Brown et al., 2020; Kaplan et al., 2020) empowers LLMs to execute target tasks by presenting specific task examples as conditions during inference, without updating model parameters. Inspired by this, ScratchpadGPT (Nye et al., 2021) enhances its proficiency in multi-step computations by mandating the model to output intermediate computation steps into a designated “scratchpad.” Codex-math (Drori et al., 2022) refines Codex by generating programs through code and integrates few-shot learning to automatically create programs for solving mathematical problems. This method has substantially improved the previous state-of-the-art accuracy in automatic solutions by 73.1% across various benchmarks like MIT’s mathematics courses, Columbia University’s Computational Linear Algebra, and the MATH benchmark (Hendrycks et al., 2021b). Notably, few-shot learning contributed to a 10% accuracy boost. Recent studies also highlight the variability in in-context learning performance across different chosen examples (Liu et al., 2022). Fu et al. (Fu et al., 2023) and Zhang et al. (Zhang et al., 2023c) selected intricate and diverse examples to enhance reasoning performance.
Figure 7. Use a “scratchpad” to improve arithmetic calculation in ScratchpadGPT (Nye et al., 2021).

### v4

@@ 2.2.2. Theorem Proving
Ongoing research is delving into the integration of LMs and interactive proof assistants within the realm of mathematics. This involves generating proofs through tactic prediction (Polu and Sutskever, 2020b; Han et al., 2022; Lample et al., 2022; Jiang et al., 2022), automating formalization processes (Wu et al., 2022; Jiang et al., 2023b), and developing unified tools (Welleck and Saha, 2023). Due to the substantial computational demands of exploration, LMs applied in this domain have traditionally been limited in size. However, recent advancements have showcased potential in employing larger models (First et al., 2023; Jiang et al., 2023b). LLEMMA (Azerbayev et al., [n. d.]) demonstrates few-shot proof autoformalization and tactic prediction, offering a vast dataset of formal mathematics and an openly accessible model for further investigation in these directions. These endeavors establish crucial theoretical and practical frameworks for automating and enhancing tactic prediction.
@@ 3. PLMs-based Methods
LMs  (§3.1) [ GPT-$f$ (Polu and Sutskever, 2020a), EPT (Kim et al., 2020), Generate &\&& Rank (Shen et al., 2021), Thor (Jiang et al., 2022), HTPS (Lample et al., [n. d.]), Galactica (Taylor et al., 2022), MATH-PLM (Hendrycks et al., 2021b), LISA (Jiang et al., 2021),
PACT (Han et al., 2022), Minerva (Lewkowycz et al., 2022), LIME (Wu et al., [n. d.]) , leaf, text width=54.2em ] ] [ Non-Autoregression
LMs  (§3.2) [ Aristo (Clark et al., 2021), GenBERT (Geva et al., 2020), NF-NSM (Feng et al., 2022), MWP-BERT (Liang et al., 2022), MathBERT(Peng et al., 2021), TAGOP (Zhu et al., 2021), MT2Net (Zhao et al., 2022),
@@
Building [ Auto-explanation (NAKAMOTO et al., 2023), ProofNet (Azerbayev et al., 2023), WizardLM (Xu et al., 2023), Wizardmath (Luo et al., 2023), ControlMath (Chen et al., 2024c),
EURUS (Yuan et al., 2024), RefAug (Zhang et al., 2024a) , leaf, text width=44.6em ] ] [ Instruct
Tuning [ MathGLM (Yang et al., 2023), Goat (Liu and Low, 2023), Calculon (Muffo et al., [n. d.]), PaLM 2-L-Math (Liu et al., 2023b), LLEMMA (Azerbayev et al., [n. d.]), Qwen-Math (Yang et al., 2024),
REFT (Luong et al., 2024), InternLM-Math (Ying et al., 2024) , leaf, text width=44.6em ] ] [ In-context
Learning [ ScratchpadGPT (Nye et al., 2021), Codex-math (Drori et al., 2022), In-Context Sampling (Liu et al., 2022), ComplexPrompt (Fu et al., 2023),
ART (Zhang et al., 2023e), LAMBADA (Kazemi et al., 2022) , LogicSolver (Yang et al., 2022) , leaf, text width=44.6em ] ] ] [ Tool-based
@@
Methods [ SymbLLM (He-Yueya et al., 2023), PoT (Chen et al., 2022a), Codex-math (Drori et al., 2022), MathPrompter (Imani et al., 2023), PAL (Gao et al., 2023), Code-scratchpad (Upadhyay et al., 2023),
MuMath-Code (Yin et al., 2024), LeanReasoner (Jiang et al., 2024), LINC (Olausson et al., 2023), , leaf, text width=44.6em ] ] [ Multi-tool
Methods [ Toolformer (Schick et al., 2023), Chameleon (Lu et al., 2023b), ART (Paranjape et al., 2023), ToolkenGPT (Hao et al., 2023b), CRITIC (Gou et al., 2023a), TALM (Parisi et al., 2022),
@@
CoT [ Answer Rationales (Ling et al., 2017), MathQA (Amini et al., 2019), CoT (Wei et al., 2022), Complexity-based CoT (Fu et al., 2023), RFT (Yuan et al., 2023a), AFT (Wang et al., 2023a),
Thought Propagation (Yu et al., 2023a), PoT (Chen et al., 2022a), AOT (Sel et al., 2023), MAmmoTH (Yue et al., 2023), SymbCoT (Xu et al., 2024a) , leaf, text width=44.6em ] ] [ Construction of
CoT [ Zero-shot CoT (Kojima et al., [n. d.]), Auto-CoT (Zhang et al., 2023e), Complexity-based CoT (Fu et al., 2023), PromptPG-CoT (Lu et al., 2023c),
AutoMate CoT (Shum et al., 2023), BoostedPrompt (Pitis et al., 2023), CoT-Influx (Huang et al., 2024) , leaf, text width=44.6em ] ] ] [ Advanced CoT
@@
Methods [ Code-based self-verification (Zhou et al., 2023d), VerifyCoT (Ling et al., 2023), DIVERSE (Li et al., 2023b), Verify-and-Edit (Zhao et al., 2023a),
Retrieval-CoT (He et al., 2022), SCREWS (Shridhar et al., 2023), SSC-CoT (Zhao et al., 2024a) , leaf, text width=44.6em ] ] [ Ensemble-based
Methods [ Self-Consistency (Wang et al., 2023d), Diversity-of-Thought (Naik et al., 2023), Complexity-based CoT (Fu et al., 2023), DIVERSE (Li et al., 2023b),
@@
Methods [ ToT (Yao et al., 2023b), TouT (Mo and Xin, 2023), ToT-system (Long, 2023), GoT (Yao et al., 2023a; Besta et al., 2023; Lei et al., 2023), RESPROMPT (Jiang et al., 2023a), RAP (Hao et al., 2023a), LATS (Zhou et al., 2023e),
LLM+P (Liu et al., 2023a), LLM+DP (Dagan et al., 2023), Self-refine (Madaan et al., 2023; Sun et al., 2023), ISR-LLM (Zhou et al., 2023b), Reflexion (Shinn et al., 2023), CR (Zhang et al., 2023c)
Step-Plan (Zhang et al., 2023b) , leaf, text width=44.6em ] ] [ Self-correct
Methods [ o1 (OpenAI, 2024), QWQ (Team, 2024), STaR (Zelikman et al., 2022), V-STaR (Hosseini et al., 2024), Quiet-STaR (Zelikman et al., 2024), REFT (Trung et al., 2024), SCoRe (Kumar et al., 2024), HGS-PRM (Ma et al., 2023)
MCTSr (Zhang et al., 2024b), CoRe (Zhu et al., 2022), Marco-o1 (Zhao et al., 2024c), PRM (Lightman et al., 2024), Step-DPO (Lai et al., 2024), Flow-DPO (Deng and Mineiro, 2024), OmegaPRM (Luo et al., 2024)
, leaf, text width=44.6em ] ] [ Socratic Teaching
Methods [ Socratic Models (Zeng et al., 2022), SOCRATIC QUESTIONING (Qi et al., 2023), Socratic prompt (Chang, 2023), SocratiQ (Al-Hossami et al., 2023b),
multi-turn Socratic advice (Ang et al., 2023), SocraticLLM (Ding et al., 2024), MATHDIAL (Macina et al., 2023) , leaf, text width=44.6em ] ] ] [ Multi-modal
Methods  (§4.5) [ Multi-modal
Reasoning [ GPT-4V (OpenAI, 2023), Qwen2-VL (Wang et al., 2024a), Gemini (Team et al., 2023), GPT-4o (Hurst et al., 2024), GLM-4V (GLM et al., 2024), LLaVA-o1 (Xu et al., 2024b), AtomThink (Xiang et al., 2024),
Math-LLaVA (Shi et al., 2024), M-STAR (Anonymous, 2024), UnAC (wang and Fu, 2024),
, leaf, text width=44.6em ] ] ] ] ]
@@
3.1. Autoregression LMs
The distinctions between Autoregression LMs and Non-Autoregressive LMs in sequence generation tasks are depicted in Figure 4. Furthermore, Autoregression LMs can be categorized into two parts: causal decoder (e.g., GPT-1 (Radford et al., 2018) and GPT-2 (Radford et al., 2019)) and encoder-decoder (e.g., T5 (Raffel et al., [n. d.])). The generation mechanisms of these two structures for sequences are depicted on the left and in the middle of Figure 4, respectively. These two architectures have not only significantly propelled the advancement of PLMs but have also become the predominant foundational architectures for LLMs.
Many improved Autoregression LMs-based methods have been proposed to enhance the performance of PLMs in mathematical problems. GPT-$f$ (Polu and Sutskever, 2020a) presents an automated prover and proof assistant to explore the performance of LMs to automated theorem proving. To resolve challenges related to expression fragmentation and operand-context separation, Kim et al. (Kim et al., 2020) introduce the Expression-Pointer Transformer (EPT), a pure neural model that employs both an ‘Expression’ token and operand-context pointers during the generation of solution equations. Generate &\&& Rank (Shen et al., 2021) introduces a novel ranking task for MWPs, a multi-task framework built upon a generative PLM, effectively addressing the challenge of minor mistakes in mathematical expressions. Thor (Jiang et al., 2022) introduces a framework that integrates language models and automated TP. The latter is employed to selectively choose relevant premises from an extensive library to assist the language model in TP. HTPS (Lample et al., [n. d.]) presents HyperTree Proof Search, a novel algorithm leveraging online training from past proof searches, enhancing model performance and surpassing the previous SOTA GPT-$f$. Galactica (Taylor et al., 2022) proposes a working memory token approach that can achieve strong performance over existing methods on mathematical MMLU (Hendrycks et al., 2021a) and MATH (Hendrycks et al., 2021b) benchmarks.
Furthermore, MATH-PLM (Hendrycks et al., 2021b) introduces a new dataset named MATH, consisting of challenging competition mathematics problems. It is found that relying solely on increasing budgets and model parameter counts would be impractical for achieving robust mathematical reasoning. LISA (Jiang et al., 2021) gathers an extensive collection of lemmas and theorems extracted from the Isabelle standard library and the Archive of Formal Proofs (AFP). Using this vast corpus, it builds LMs that prove theorems effectively within the AFP. PACT (Han et al., 2022) extracts abundant self-supervised data from kernel-level proof terms, enabling joint training alongside the standard tactic prediction objective. This effectively enhances the model’s ability for TP. Minerva (Lewkowycz et al., 2022) trains a language model, initially pre-trained on general natural language data, through additional training on technical content to address Quantitative Reasoning Problems. It achieves a SOTA performance on technical benchmarks including MATH (Hendrycks et al., 2021b). LIME (Wu et al., [n. d.]) introduces a novel pre-training methodology that specifically learns inductive bias for mathematical reasoning. Observations indicate that models trained using LIME exhibit superior performance compared to standard transformers.
@@ 3.2. Non-Autoregression LMs
Unlike Autoregression LMs, Non-Autoregression LMs enable the model to generate all parts of a sequence simultaneously, without depending on previously generated content, as depicted on the right side of Figure 4. Notably, models like BERT (Kenton and Toutanova, 2019), Roberta (Liu et al., 2019), and related architectures utilize masked word representations, leveraging pre-trained context-aware embeddings as comprehensive semantic features. This significantly elevates the performance benchmarks of NLP tasks.
@@ 4. LLMs-based Methods
Large-scale language models (LLMs) are designed for processing and generating text akin to human communication (Touvron et al., 2023a; OpenAI, 2023). Mathematics is also a form of language, that communicates complex concepts and relationships through a structured system of symbols and notation, akin to the rules of a spoken language. Thus, a language model that grasps these mathematical rules can ”speak” the language of mathematics, proving to be a valuable asset for mathematicians. In numerous ways, a sophisticated language model like GPT-4 (OpenAI, 2023) becomes an invaluable tool in the field of mathematics (Azerbayev et al., [n. d.]; Luo et al., 2023). We classify the existing studies into four parts: instruction learning, tool-based methods, fundamental CoT methods, and advanced CoT methods (Figure 3).
@@ 4.1. Instruction Learning
Instruction Building refers to the process of designing and generating high-quality instruction data that guide models in improving specific capabilities, such as mathematical reasoning. After pre-training, these instructions are crucial for directing models towards more sophisticated understanding and application of mathematical concepts. A semi-supervised approach, as presented in Auto-explanation (NAKAMOTO et al., 2023), leverages LLMs to create datasets for automating the scoring of mathematical self-explanations in the context of mathematics education. Similarly, RefAug (Zhang et al., 2024a) leverages GPT-4o to construct reflective augmentation data for training model, thereby enhancing their performance on mathematical reasoning tasks. A challenging benchmark ProofNet (Azerbayev et al., 2023) is introduced to build a system of automated theorem proving. Meanwhile, prompt retrieval and distilled back translation methods are introduced for statement autoformalization. WizardLM (Xu et al., 2023) presents a groundbreaking method called Evol-Instruct, designed to autonomously generate high-quality instructions by LLMs themselves. Furthermore, Wizardmath (Luo et al., 2023) proposes a reinforced Evol-instruct Feedback method (RLEIF) to build more complexity of instruct dataset. This method effectively enhances the mathematical reasoning capabilities of Llama-2 by supervised fine-tuning and –Proximal Policy Optimization (PPO) training (Schulman et al., 2017). ControlMath (Chen et al., 2024c) introduces a novel approach that first generates diverse equations, which are then transformed into math word problems by a Problem Crafter LLM-agent. EURUS (Yuan et al., 2024) generates a large number of reasoning trajectories by leveraging CoT (Wei et al., 2022) and tool creation (Qian et al., 2023) methods. These trajectories within multi-turn interactions (Wang et al., 2023c) are organized into a preference tree structure. Then extensive instruction and preference data are collected from this tree to enhance the model’s mathematical reasoning capabilities by Direct Preference Optimization (DPO) (Rafailov et al., 2024), Kahneman-Tversky Optimization (KTO) (Ethayarajh et al., 2024), and Noise Contrastive Alignment (NCA) (Chen et al., 2024a).
@@
Instruction tuning stands as a potent method to elevate the prowess of large models, aiding in steering the model towards generating outputs that align with human intent (Zhang et al., 2023a). Compared to the pre-training phase of LLMs, instruction tuning demands notably less data and computational resources, making it a prevalent strategy for enhancing a model’s domain-specific capabilities. For instance, MathGLM (Yang et al., 2023) showcases that even an LLM with 2B parameters can adeptly execute multi-digit arithmetic operations with limited training data. Goat (Liu and Low, 2023) discerns patterns in the tokenization of different numbers and proposes a consistent approach for their tokenization based on the LLaMA model. Similarly, Calculon (Muffo et al., [n. d.]) refines the model’s arithmetic skills by employing a digit decomposition technique to construct a fine-tuning dataset. In PaLM 2-L-Math (Liu et al., 2023b), three fine-tuning strategies were explored, revealing key insights: 1) The quality and style of step-by-step solutions significantly influence model performance; 2) Combining solution re-ranking and majority voting yields a more substantial performance enhancement compared to using them individually. Additionally, multi-task fine-tuning, separating solution generation and evaluation tasks, surpasses the baseline achieved by solution fine-tuning alone. Moving beyond the enhancement of mathematical capabilities through fine-tuning, LLEMMA (Azerbayev et al., [n. d.]) introduces the Proof-Pile-2 dataset, an amalgamation of mathematical texts and code. Continual pre-training with Code Llama empowers the model to leverage Python interpreters and formal theorem provers, demonstrating impressive performance on the MATH benchmark. To improve the effectiveness of instruction tuning, Luong et al. (Luong et al., 2024) introduce an approach that incorporates online reinforcement learning. This method enhances the model’s generalization by exploring multiple CoTs and using real-world answers as reward signals. Recently, more math-specific LLMs (Yang et al., 2024; Ying et al., 2024) have been proposed, with enhanced instruction tuning tailored for the mathematical domain. Qwen2.5-MATH (Yang et al., 2024) begins by pretraining on large-scale, high-quality mathematical data. In the post-training phase, it develops a reward model (RM), which is then utilized to the iterative evolution of data in supervised fine-tuning (SFT). Thereafter, both the SFT model and the RM undergo iterative training. During the inference stage, the RM guides the sampling process to optimize the model’s performance. In contrast, InternLM-Math (Ying et al., 2024) undergoes secondary pretraining on a base model using a mathematical corpus. In the post-training phase, it is fine-tuned with instructions through unifying CoT and code interpretation, reward modeling, data augmentation under an unified seq2seq format.
@@
In-context Learning (ICL) (Brown et al., 2020; Kaplan et al., 2020) empowers LLMs to execute target tasks by presenting specific task examples as conditions during inference, without updating model parameters. Inspired by this, ScratchpadGPT (Nye et al., 2021) enhances its proficiency in multi-step computations by mandating the model to output intermediate computation steps into a designated “scratchpad.” Codex-math (Drori et al., 2022) refines Codex by generating programs through code and integrates few-shot learning to automatically create programs for solving MWPs. This method has substantially improved the previous state-of-the-art accuracy in automatic solutions by 73.1% across various benchmarks like MIT’s mathematics courses, Columbia University’s Computational Linear Algebra, and the MATH benchmark (Hendrycks et al., 2021b). Notably, few-shot learning contributed to a 10% accuracy boost. Recent studies also highlight the variability in in-context learning performance across different chosen examples (Liu et al., 2022). Fu et al. (Fu et al., 2023) and Zhang et al. (Zhang et al., 2023e) selected intricate and diverse examples to enhance reasoning performance. LAMBADA (Kazemi et al., 2022) introduces a backward reasoning method for mathematical inference, starting from conclusions and working back to axioms. It uses few-shot to guide LLMs in reverse problem-solving, achieving promising results. LogicSolver (Yang et al., 2022) first constructed a dataset of mathematical word problems, InterMWP, where each problem includes interpretable logical formulas based on algebraic knowledge. The relevant logical formulas in InterMWP is used as In-context to enhance both interpretability and accuracy.